VIBRATIONAL COUPLING PATHWAYS IN THE CH STRETCH REGION OF CH_{3}OH AND CH_{3}OD AS REVEALED BY IR AND FTMW-IR SPECTROSCOPIES

\underline{SYLVESTRE TWAGIRAYEZU}, XIAOLIANG WANG, AND DAVID S. PERRY, Department of Chemistry, The University of Akron, Akron OH 44325; JUSTIN L. NEILL, MATT T. MUCKLE, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904; LI-HONG XU, Department of Physics, Centre for Laser, Atomic and Molecular Studies (CLAMS), University of New Brunswick, Saint John, New Brunswick E2L 4L5, Canada.

Infrared spectra of jet-cooled CH_{3}OD and CH_{3}OH in the CH stretch region are observed by coherence-converted population transfer Fourier transform microwave-infrared (CCPT-FTMW-IR) spectroscopy (E torsional species only) and by slit-jet single resonance spectroscopy (both A and E torsional species, CH_{3}OH only). Previously, we reported^{a} the analysis of ν_{3} symmetric CH stretch region (2750−2900 cm^{−1}), and the present work extends the analysis to higher frequency (2900−3020 cm^{−1}). The overall observed spectra contain 17 interacting vibrational bands for CH_{3}OD and 28 for CH_{3}OH. The signs and magnitudes of the torsional tunneling splittings are deduced for three CH fundamentals (ν_{3}, ν_{9}, ν_{2}) of both molecules and are compared to a model calculation and to ab initio theory. The number and distribution of observed vibrational bands indicate that the CH stretch bright states couple first to doorway states that are binary combinations of bending modes. In the parts of the spectrum where doorway states are present, the observed density of coupled states is comparable to the total density of vibrational states in the molecule, but where there are no doorway states, only the CH stretch fundamentals are observed. A time-dependent interpretation of the present FTMW-IR spectra indicates a fast (∼ 200 fs) initial decay of the bright state followed by second, slower redistribution (∼ 1−3 ps). The qualitative agreement of the present data with the time-dependent experiments of Iwaki and Dlott ^{b} provides further support for the similarity of the fastest vibrational relaxation processes in the liquid and gas phases.

aTwagirayezu, S.; Clasp, T. N.; Perry, D. S.; Neill, J. L.; Muckle, M. T.; Pate, B. H. J. Phys. Chem. A 2010, 114, 6818

bIwaki, L. K.; Dlott, D. D. J. Phys. Chem. A 2000, 104, 9101